### v1

# InvZW: Invariant Feature Learning via Noise-Adversarial Training for Robust Image Zero-Watermarking

Abstract
This paper introduces a novel deep learning framework for robust image zero-watermarking based on distortion-invariant feature learning. As a zero-watermarking scheme, our method leaves the original image unaltered and learns a reference signature through optimization in the feature space. The proposed framework consists of two key modules. In the first module, a feature extractor is trained via noise-adversarial learning to generate representations that are both invariant to distortions and semantically expressive. This is achieved by combining adversarial supervision against a distortion discriminator and a reconstruction constraint to retain image content. In the second module, we design a learning-based multibit zero-watermarking scheme where the trained invariant features are projected onto a set of trainable reference codes optimized to match a target binary message. Extensive experiments on diverse image datasets and a wide range of distortions show that our method achieves state-of-the-art robustness in both feature stability and watermark recovery. Comparative evaluations against existing self-supervised and deep watermarking techniques further highlight the superiority of our framework in generalization and robustness.
Index Terms:
Robust Zero-watermarking, invariant feature learning, adversarial training, representation learning, deep watermarking
I Introduction
Image watermarking is a widely used technique for embedding hidden information into digital media for purposes such as copyright protection and content authentication. In contrast to standard watermarking methods that modify the cover image by embedding a watermark directly into its pixel or transform domain representation [1], zero-watermarking leaves the original image unchanged. Instead, it extracts robust features from the image and associates them with a watermark to form a signature, which is stored externally without altering the image. During usage, features extracted from a potentially marked image are used to either verify the watermark or recover it. Zero-watermarking is particularly useful in domains where modifying the original image is undesirable, such as in medical diagnostics, legal evidence, and satellite imaging. There are two primary objectives of zero-watermarking: verification-based approaches, which confirm the presence of a watermark [2], and multibit extraction-based methods, which aim to retrieve the full watermark content from the extracted features [3].
Recent advancements in deep learning have transformed the landscape of zero-watermarking by introducing learning-based approaches that leverage deep neural networks to extract high-level feature representations from images [4]. Instead of relying on handcrafted descriptors, these methods employ convolutional networks or transformers to learn features that are more robust to distortions [3, 5, 6]. A common approach involves using a deep network as a fixed feature extractor: features are drawn from the input image and combined with a watermark to generate a reference signature. During verification, the same feature extraction process is applied to a test image to determine watermark presence or recover watermark bits (see Fig. 1). Beyond zero-watermarking, several deep image watermarking schemes [7, 8] also leverage pre-trained networks, especially self-supervised models, as fixed feature extractors. These methods modify the input image to embed a watermark that becomes detectable in the feature space. By inheriting the robustness of self-supervised representations, they offer robustness to distortions without additional training. These approaches typically assume that off-the-shelf models, trained on large-scale image classification or self-supervised learning tasks, provide sufficient invariance to potential image distortions in watermarking.
Figure 1: Image Zero-Watermarking with Deep Learning.
Robustness is a critical consideration in image zero-watermarking, where the extracted features must remain stable under a wide range of distortions. These include common image processing attacks such as blurring and compression, as well as geometric transformations such as rotation, scaling, and cropping. These transformations can significantly alter image appearance while preserving semantics, posing a challenge for watermark retrieval. While existing methods have leveraged pre-trained feature extractors to achieve a degree of robustness, they are not explicitly designed for watermarking and often fall short when encountering unseen or complex distortions. This highlights the need for more specialized approaches that directly target invariance in the feature space, tailored to the demands of watermarking applications.
To this end, we propose a noise-adversarial training framework for learning invariant features specifically designed for multibit image zero-watermarking. Our method introduces three core novelties: (1) a new invariant feature learning scheme that is explicitly tailored for watermarking applications, ensuring that feature representations remain stable under diverse distortions; (2) a novel deep learning-based zero-watermarking method that leverages these learned features for watermark extraction without modifying the original image; and (3) a significant improvement in robustness, as our adversarial training approach enables the model to withstand a wide range of image processing and geometric transformations. Collectively, these components constitute a framework that advances the robustness and reliability of deep zero-watermarking.
II Related Work
This section reviews the most relevant areas, including invariant feature learning and deep learning-based zero-watermarking.
II-A Invariant Feature Learning
Robust feature learning has its origins in convolutional neural networks (CNNs), which introduce some invariance and equivariance through architectural components such as convolutional filters and pooling operations [9]. Recent advancements have shifted towards unsupervised and data-driven approaches for determining robustness, reflecting a significant change in paradigm [10, 11, 12]. A key development is the use of autonomous invariance learning through methods like differentiable Kronecker-factored Laplace approximations, which eliminate the need for manual data augmentation selection by optimizing augmentation parameters directly during training [13]. This enables models to learn transformation robustness from data, adapting to varied conditions without predefined assumptions. While Bayesian approaches to invariance learning have been successfully applied in simpler models like Gaussian processes [14], extending these methods to deep networks remains a challenge. Some advances in scalable marginal likelihood approximations, such as those using the structured Gauss-Newton method, have made it feasible to apply these techniques in deep learning [15, 16]. This evolution marks a shift from predefined augmentation strategies towards adaptive, data-driven methods, which align closely with the goals of robust feature extraction.
Self-supervised learning (SSL) methods, such as SimCLR [17], BYOL [18], and DINO [19], contribute to this trend by implicitly learning invariance through contrastive objectives or teacher-student frameworks during training. However, these methods primarily focus on improving the quality of learned representations, not explicitly on enhancing robustness against a wide range of distortions. Some work has attempted to address this gap in graph neural networks through methods, which employs tailored objectives to disentangle invariant features from spurious correlations [20].
Our approach extends these principles by directly training a deep learning model for invariance, with a specific focus on enhancing robustness. Unlike SSL approaches that treat invariance as a byproduct, our method is designed to optimize feature stability under diverse image transformations. This makes it particularly suitable for applications of image watermarking, where robustness to distortions is critical.
II-B Deep Learning-based Image Zero-Watermarking
A common strategy behind recent deep learning-based image zero-watermarking methods is to adopt an existing deep model, typically pre-trained on large-scale datasets, as a feature extractor and assume that the learned features offer sufficient robustness for watermarking tasks. Fierro et al. [3] utilized a CNN to extract features from the cover host image, which were then combined with a permuted binary watermark using an XOR operation to generate a master share for later verification. He et al. [5] extended this pipeline by introducing fully connected layers to extract shallow features from multiple convolutional layers, while incorporating a shrinkage module for soft thresholding and a noise layer during training to enhance robustness. Han et al. [6] further improved this approach by encrypting the watermark using a chaotic encryption scheme and employing the Swin Transformer [21] as the feature extractor to improve geometric invariance and overall robustness. Li et al. [2] proposed a zero-watermarking framework that uses ConvNeXt and LK-PAN to extract local and global features, with a dedicated watermark block designed to enhance both robustness and discriminability without handcrafted features.
Beyond zero-watermarking, several deep image watermarking schemes [7, 8] have explored the use of pre-trained networks—particularly self-supervised models—as fixed feature extractors. These methods embed watermarks by optimizing the input image while keeping the model weights frozen, allowing the watermark to be revealed in the feature space. By leveraging the inherent robustness of self-supervised representations, these approaches achieve resilience to common distortions without the need for retraining the feature extractor.
Current approaches typically assume that off-the-shelf models, trained for image classification or self-supervised tasks, offer sufficient invariance for watermarking. In contrast, we train a deep network specifically to learn invariant features tailored for watermarking, resulting in greater robustness across diverse transformations. Moreover, unlike many existing methods that rely on manually computing feature-watermark correlations for verification, our approach introduces a novel learning-based mechanism that derives the reference signature from the extracted invariant features and the target watermark, enabling more reliable and adaptive zero-watermarking.
III Proposed Method
To simplify the discussion, we conceptually decompose the proposed framework into two primary modules, as illustrated in Fig. 2. Module 1 performs invariant feature learning, where a deep network is trained to extract robust image features that remain stable under diverse distortions. Module 2 introduces a learning-based zero-watermarking scheme, which combines the extracted features with a given watermark to generate a reference signature for later extraction. The remainder of this section discusses the details of each module.
Figure 2: Modules of the Proposed Method.
III-A Module 1: Invariant Feature Learning via Noise-Adversarial Training
Fig. 3 illustrates the proposed invariant feature learning module based on noise-adversarial training. The framework consists of three main components: a feature extractor, a discriminator, and a reconstructor. Given both original and distorted images, the feature extractor—implemented using a Transformer-based architecture—generates high-level representations intended to be invariant to distortions. The discriminator is trained to classify whether a given feature originates from an original or distorted image, while the feature extractor is trained adversarially to fool the discriminator. This minimax interaction encourages the extractor to produce features that are indistinguishable across distortions, thereby promoting invariance. To ensure that the learned features retain semantic content, the reconstructor is trained to reconstruct the original image from these features, forcing the invariant features to preserve meaningful information relevant to the image itself.
Figure 3: Module 1: Proposed Invariant Feature Learning via Noise-Adversarial Training.
III-A1 Feature Extractor
The feature extractor $\mathrm{FE}$ is implemented using the Vision Transformer (ViT) architecture [22], which is adapted here to learn features that are robust to image distortions. Given an input image $x\in\mathbb{R}^{128\times 128\times 3}$, the image is divided into $N=64$ non-overlapping patches, each of size $16\times 16$. Each patch $p_{i}$ is flattened and projected into a $d$-dimensional latent space via a learnable linear transformation: $z_{i}=W_{p}p_{i}+b_{p}$, where $W_{p}\in\mathbb{R}^{d\times(16\times 16\times 3)}$ and $b_{p}\in\mathbb{R}^{d}$. To encode positional information, learnable positional encodings $E_{\text{pos},i}$ are added to each patch embedding, resulting in $z_{i}^{(0)}=z_{i}+E_{\text{pos},i}$.
The sequence of embedded patches is then passed through $L=12$ stacked Transformer blocks, each consisting of a multi-head self-attention (MHSA) layer followed by a feed-forward network (FFN). Formally, each block applies $z^{(\ell)}=\mathrm{LayerNorm}(z^{(\ell-1)}+\mathrm{MHSA}(z^{(\ell-1)}))$, followed by $z^{(\ell)}=\mathrm{LayerNorm}(z^{(\ell)}+\mathrm{FFN}(z^{(\ell)}))$, for $\ell=1,\dots,L$.
From the final Transformer block, we extract the output corresponding to the [CLS] token as the global representation of the image, denoted as $F=\mathrm{FE}(x)=z^{(L)}_{\text{CLS}}$. Both the original and distorted images are passed through the same feature extractor, and the model will be trained using adversarial objectives to ensure that the extracted features remain invariant to distortions. These features are then passed to the discriminator and reconstructor for further supervision, as detailed in the following subsections.
III-A2 Discriminator
The discriminator is designed to classify whether a feature representation originates from an original or a distorted image. It follows a ViT-based architecture, structurally similar to the feature extractor, but adapted for binary classification.
Given an input image of size $128\times 128\times 3$, the discriminator first divides it into 64 non-overlapping patches of size $16\times 16$. Each patch is flattened and linearly projected into a 64-dimensional latent space using a learnable embedding layer. To preserve spatial ordering, positional embeddings are added to each projected patch vector. A learnable [CLS] token is prepended to the patch sequence, serving as a global aggregation token for classification.
The sequence, including the [CLS] token, is passed through a stack of $L=8$ Transformer blocks. Each block contains a MHSA layer followed by a FFN as in Fig. 3, with residual connections and layer normalization applied at each submodule. This architecture allows the model to capture both local and global relationships across the image patches.
After Transformer processing, the output corresponding to the [CLS] token is extracted and passed through a dense classification head consisting of a single linear layer followed by a sigmoid activation: $D(F)=\sigma(W_{d}\cdot F_{\text{CLS}}+b_{d}),$ where $F_{\text{CLS}}$ is the final embedding of the classification token, $W_{d}\in\mathbb{R}^{1\times d}$ and $b_{d}\in\mathbb{R}$ are the learnable weight and bias parameters of the classification head, and $\sigma$ denotes the sigmoid function.
The discriminator will be trained to distinguish between features derived from original and distorted images. Its gradients provide adversarial feedback to the feature extractor, encouraging the extractor to produce representations that are invariant with respect to the applied distortions.
III-A3 Reconstructor
The reconstructor, illustrated in Fig. 4, can be conceptually described as an autoencoder-like architecture designed to enforce semantic preservation in the learned features. It consists of an encoder path, a bottleneck, and a decoder path. The input to the reconstructor is the feature tensor $F\in\mathbb{R}^{128\times 128\times 3}$ produced by the feature extractor. This reconstruction objective ensures that the extracted features retain sufficient semantic information, discouraging the feature extractor from collapsing into trivial or non-informative representations.
Figure 4: Proposed Reconstructor Architecture.
The encoder comprises a series of convolutional blocks, each consisting of a 2D convolution layer followed by batch normalization and ReLU activation. These layers progressively reduce the spatial resolution from $128\times 128$ to $8\times 8$, while increasing the channel depth from 3 to 512. A 2D max pooling operation is applied at each stage to downsample the spatial dimensions. The resulting compressed representation at the bottleneck has dimensions $8\times 8\times 512$, capturing high-level semantic information in a compact form. The decoder mirrors the encoder structure, utilizing transposed convolution layers to upsample the feature maps and gradually restore the spatial resolution. The channel dimensions are symmetrically reduced to reconstruct the image in its original shape. The final output of the decoder is a reconstructed image $\hat{I}=f_{\text{recon}}(F)$, where a sigmoid activation is applied in the last layer to constrain pixel values between 0 and 1.
III-A4 Training
The objective of the training process is to learn a feature extractor $\mathrm{FE}$ that produces representations invariant to image distortions, while preserving semantic information necessary for downstream tasks such as watermarking. The training consists of three components: updating the discriminator $D$, the feature extractor $\mathrm{FE}$, and the reconstructor $R$. Given an input image $x\sim p(x)$, and its distorted version $x^{\prime}\sim p_{T}(x)$, we define the following loss functions and update routines.
Training the Discriminator. The discriminator $D$ is trained to distinguish between features extracted from clean and distorted images. It is optimized using a binary cross-entropy loss:
$\mathcal{L}_{D}=$ $\;\mathbb{E}_{x\sim p(x)}\left[\log D(\mathrm{FE}(x))\right]$
$+\mathbb{E}_{x^{\prime}\sim p_{T}(x)}\left[\log\left(1-D(\mathrm{FE}(x^{\prime}))\right)\right].$ (1)
This objective encourages $D$ to assign high confidence to features extracted from clean images and low confidence to those from distorted images.
Training the Feature Extractor. The feature extractor is trained adversarially to fool the discriminator, making the distributions of $\mathrm{FE}(x)$ and $\mathrm{FE}(x^{\prime})$ indistinguishable. The adversarial loss is given by:
$\mathcal{L}_{\mathrm{adv}}=\mathbb{E}_{x^{\prime}\sim p_{T}(x)}\left[\log\left(1-D(\mathrm{FE}(x^{\prime}))\right)\right].$ (2)
To ensure that the features retain semantic information, the feature extractor is also optimized with a reconstruction loss via the reconstructor $R$, which reconstructs the input from the features: $\hat{x}=R(\mathrm{FE}(x)),\quad\hat{x}^{\prime}=R(\mathrm{FE}(x^{\prime})).$ The reconstruction loss is a weighted combination of the Mean Squared Error (MSE) and the Structural Similarity Index (SSIM):
$\mathcal{L}_{R}=\lambda_{S}\left(1-\mathrm{SSIM}(x,\hat{x})\right)+\lambda_{M}\|x-\hat{x}\|_{2}^{2}.$ (3)
This loss is applied to both clean and distorted inputs, enforcing that features from either domain encode sufficient information for image recovery. The overall objective for the feature extractor combines the adversarial and reconstruction losses:
$\mathcal{L}_{\mathrm{FE}}=\lambda_{D}\cdot\mathcal{L}_{\mathrm{adv}}+\mathcal{L}_{R},$ (4)
where $\lambda_{D}$ controls the strength of the adversarial term. By minimizing $\mathcal{L}_{\mathrm{FE}}$, the feature extractor learns to produce invariant but semantically rich representations.
Training the Reconstructor. The reconstructor $R$ is trained jointly with $\mathrm{FE}$ to reconstruct the original image from the learned features. The same loss $\mathcal{L}_{R}$ as in Eq. (3) is minimized with respect to $\theta_{R}$, the parameters of the reconstructor. This supervision ensures that the features do not collapse to trivial encodings and remain informative. Algorithm 1 describes the proposed adversarial training for invariant feature learning.
Algorithm 1 Noise-Adversarial Training for Invariant Feature Learning
0:  Feature Extractor $\mathrm{FE}$, Discriminator $D$, Reconstructor $R$, transformation function $T$, dataset $X$, batch size $B$, learning rate $\alpha$, loss weights $\lambda_{1},\lambda_{2}$.
1:  for each epoch do
2:     for each batch of images $x$ sampled from $X$ do
3:        Generate distorted images: $x^{\prime}=T(x)$.
4:        Extract features: $F=\mathrm{FE}(x)$, $F^{\prime}=\mathrm{FE}(x^{\prime})$.
5:        Train Discriminator:
6:        Compute labels: $y=1$ for $F$, $y=0$ for $F^{\prime}$.
7:        Compute discriminator loss $\mathcal{L}_{D}$.
8:        Update $D$ using $\theta_{D}\leftarrow\theta_{D}-\alpha\nabla_{\theta_{D}}\mathcal{L}_{D}$.
9:        Train Feature Extractor:
10:        Compute adversarial loss $\mathcal{L}_{\mathrm{adv}}$.
11:        Compute reconstruction loss $\mathcal{L}_{R}$.
12:        Compute total loss $\mathcal{L}_{\mathrm{FE}}=\lambda_{1}\mathcal{L}_{\mathrm{adv}}+\lambda_{2}\mathcal{L}_{R}$.
13:        Update $\mathrm{FE}$ using $\theta_{\mathrm{FE}}\leftarrow\theta_{\mathrm{FE}}-\alpha\nabla_{\theta_{\mathrm{FE}}}\mathcal{L}_{\mathrm{FE}}$.
14:        Train Reconstructor:
15:        Reconstruct images: $\hat{x}=R(F)$.
16:        Compute reconstruction loss $\mathcal{L}_{R}$.
17:        Update $R$ using $\theta_{R}\leftarrow\theta_{R}-\alpha\nabla_{\theta_{R}}\mathcal{L}_{R}$.
18:     end for
19:     Save model weights every 10 epochs.
20:  end for
III-A5 Theoretical Justification
The proposed noise-adversarial training framework is theoretically supported by two complementary perspectives: the Information Bottleneck (IB) principle [23] and domain adversarial learning [24].
First, we interpret the dual-objective formulation—adversarial training for invariance and reconstruction for semantic preservation—through the lens of the Information Bottleneck principle. The goal is to learn a feature representation $F$ that maximizes mutual information with the input image $x$, while minimizing mutual information with irrelevant nuisance factors introduced by distortion $x^{\prime}$. This trade-off can be formalized as:
$\max_{\mathrm{FE}}\,I(F,x)-\beta I(F,x^{\prime}),$ (5)
where $I(\cdot,\cdot)$ denotes mutual information, and $\beta>0$ balances informativeness against invariance. While mutual information is not computed explicitly, the adversarial loss serves as a surrogate to minimize $I(F,x^{\prime})$ by aligning the feature distributions of $x$ and $x^{\prime}$, i.e., $p(F)\approx p(F^{\prime})$. Simultaneously, the reconstruction loss enforces that $F$ retains sufficient information to recover the original content, implicitly maximizing $I(F,x)$.
This framework can also be viewed as a case of domain adversarial learning, where the distortion acts as a latent domain shift. In this view, the original image $x$ and its distorted version $x^{\prime}$ belong to two domains with shared semantic content but domain-specific variations (i.e., noise). The adversarial training objective ensures that the learned features are indistinguishable across domains:
$\mathbb{E}_{x\sim p(x)}\left[D(\mathrm{FE}(x))\right]=\mathbb{E}_{x^{\prime}\sim p_{T}(x)}\left[D(\mathrm{FE}(x^{\prime}))\right]=0.5.$ (6)
This alignment encourages the learned representation $F$ to be domain-invariant (i.e., distortion-invariant), such that noise does not affect the feature encoding. However, domain invariance alone is insufficient, as it may lead to degenerate solutions where the extracted features collapse to trivial or constant vectors. To prevent such feature collapse, the reconstructor imposes a functional constraint: $x\approx R(\mathrm{FE}(x)),\quad x^{\prime}\approx R(\mathrm{FE}(x^{\prime})),$ which enforces that the features remain semantically meaningful and non-trivial. Together, these constraints ensure that the feature extractor learns robust, invariant, and expressive representations—properties that are critical for downstream watermark retrieval in zero-watermarking applications.
III-B Module 2: Learning-based Zero-Watermarking with Invariant Feature
In the second module, we present a novel multibit zero-watermarking scheme that leverages the invariant feature representation learned in Module 1. As a zero-watermarking scheme, our method does not modify the original image. Instead, a reference signature is optimized per image-watermark pair such that the semantic content captured by the invariant features can be reliably associated with a binary watermark message.
III-B1 Feature Extraction and Encoding
Given an image $x$, its invariant feature representation $F=\mathrm{FE}(x)\in\mathbb{R}^{H\times W\times C}$ is extracted using the trained feature extractor from Module 1. To reduce the dimensionality and obtain a compact vector form, we apply a transformation function $\Psi$, composed of global average pooling followed by a fully connected projection, yielding:
$\tilde{F}=\Psi(F),\quad\tilde{F}\in\mathbb{R}^{d}.$ (7)
Let $W\in\{0,1\}^{k}$ be the target binary watermark message of length $k$, and let $C\in\mathbb{R}^{k\times d}$ denote a learnable reference signature matrix. The goal is to optimize $C$ such that the inner product between each watermark bit vector $C_{i}\in\mathbb{R}^{d}$ and the feature vector $\tilde{F}$ aligns with the corresponding bit $W_{i}$. The predicted watermark is computed as:
$\hat{W}_{i}=\sigma(\tilde{F}\cdot C_{i}),\quad i=1,\dots,k,$ (8)
where $\sigma(\cdot)$ is the sigmoid activation function.
III-B2 Loss Function and Optimization
To enforce alignment between the predicted watermark $\hat{W}$ and the target message $W$, we minimize a binary cross-entropy loss:
$\mathcal{L}_{W}=-\sum_{i=1}^{k}\left[W_{i}\log\hat{W}_{i}+(1-W_{i})\log(1-\hat{W}_{i})\right].$ (9)
To prevent overfitting and numerical instability, we regularize the reference signature matrix using an $L_{2}$ penalty: $\mathcal{L}_{C}=\|C\|_{2}^{2}.$ The overall optimization objective becomes: $\mathcal{L}_{\text{total}}=\mathcal{L}_{W}+\lambda_{C}\mathcal{L}_{C},$ where $\lambda_{C}$ controls the regularization strength.
The reference signature $C$ is optimized using gradient descent:
$C\leftarrow C-\eta\cdot\frac{\partial\mathcal{L}_{\text{total}}}{\partial C},$ (10)
where $\eta$ is the learning rate. This optimization is performed per image-watermark pair during registration and does not involve updating the image or feature extractor. Similarly, the gradient of the total loss $\mathcal{L}_{\text{total}}$ with respect to the parameters of the fully connected projection in $\Psi$ is used to update the weights of this layer. This allows the transformation function $\Psi$ to adapt to the watermark decoding task while keeping the feature extractor fixed.
III-B3 Watermark Extraction
To extract the watermark, the distorted image is passed through the trained feature extractor and transformation function to yield $\tilde{F}^{\prime}=\Psi(\mathrm{FE}(x^{\prime}))$. The predicted watermark is then computed as:
$\hat{W}_{i}=\sigma(\tilde{F}^{\prime}\cdot C_{i}),\quad i=1,\dots,k.$ (11)
A thresholding step is applied to recover the binary message:
$\tilde{W}_{i}=\begin{cases}1,&\hat{W}_{i}\geq 0.5,\\ 0,&\hat{W}_{i}<0.5.\end{cases}$ (12)
Robust watermark recovery is enabled by the invariance of $\mathrm{FE}$, which ensures that $\tilde{F}^{\prime}\approx\tilde{F}$ even under substantial image distortions. Since the reference signature is optimized in feature space rather than image space, and the image is left unmodified, this approach fully conforms to the principles of zero-watermarking.
III-B4 Method Advantages
A key practical advantage of the proposed formulation is its flexibility in associating different watermark messages with the same image. Since the reference signature $C$ is optimized in feature space and not embedded in the image, it can be re-trained per user or application without modifying the image or the invariant feature extractor. This supports a one-to-many mapping: a single image can be linked to multiple watermark identities, each identified by a unique $C$. During extraction, only the correct reference signature will yield a successful extraction, providing a mechanism for ownership differentiation, user-specific claims, or multi-level authorization.
From a theoretical perspective, the proposed approach can be interpreted as solving a constrained optimization problem in the invariant feature space. Given a fixed, distortion-invariant representation $\tilde{F}$, the training of $C$ seeks to project $\tilde{F}$ onto a set of learned directional codes $\{C_{1},\dots,C_{k}\}$ such that their inner products reflect a target binary vector $W$. The sigmoid activation and binary cross-entropy loss induce a soft margin between watermark bits, effectively identifying a hyperplane that separates $\tilde{F}\cdot C_{i}\geq 0.5$ from $<0.5$. By minimizing this loss, the solution converges to a directionally optimal configuration of $C$, tailored to the unique geometry of the input feature. This learning-based formulation generalizes well to unseen transformations, as the optimization finds reference codes that are not just memorized, but robustly aligned with the semantic content captured by $\tilde{F}$.
Algorithm 2 Zero Watermarking and Extraction using Invariant Features
1:  Input: Original images $\{I_{i}\}_{i=1}^{N}$, Watermarks $\{W_{i}\}_{i=1}^{N}$, Reference $C\in\mathbb{R}^{k\times d}$, Feature Extractor $FE$
2:  Output: Optimized Reference $C\in\mathbb{R}^{k\times d}$ and $\Psi$
3:  for each image $I_{i}$ do
4:     Extract invariant feature map: $F=FE(I_{i})\in\mathbb{R}^{128\times 128\times 3}$
5:     Compute $\tilde{F}=\Psi(F),\quad\tilde{F}\in\mathbb{R}^{d}$
6:     for each bit $b=1$ to $B$ do
7:        Compute dot product: $s_{i,b}=\tilde{F}_{i}\cdot C[i,b,:]$
8:     end for
9:     Apply sigmoid: $\hat{W}_{i}=\sigma(s_{i})\in\mathbb{R}^{B}$
10:     Compute Binary Cross Entropy loss: $\mathcal{L}_{W}=\text{BCE}(\hat{W}_{i},W_{i})$
11:     Regularize Reference: $\mathcal{L}_{C}=\|C\|_{2}^{2}$
12:     Total Loss: $\mathcal{L}_{total}$= $\mathcal{L}_{W}+\mathcal{L}_{C}$
13:     Update $C$ using $\theta_{C}\leftarrow\theta_{C}-\alpha\nabla_{\theta_{C}}\mathcal{L}_{total}$.
14:     Update $\Psi$ using $\theta_{\Psi}\leftarrow\theta_{\Psi}-\alpha\nabla_{\theta_{\Psi}}\mathcal{L}_{total}$.
15:  end for
16:  Watermark Extraction:
17:  Input: Noisy image $I^{\prime}$, Feature Extractor $FE$, Optimized Reference $C$, Optimized $\Psi$
18:  Extract invariant feature map: $F=FE(I_{i})$
19:  Apply $\Psi$: $\tilde{F}=\Psi(F)$
20:  for each bit $b=1$ to $B$ do
21:     Compute dot product: $s^{\prime}_{b}=\tilde{F}\cdot C[*,b,:]$
22:     Apply sigmoid: $\hat{W}_{b}=\sigma(s^{\prime}_{b})$
23:     Assign bit: $W_{b}=\begin{cases}1&\text{if }\hat{w}_{b}\geq 0.5\\ 0&\text{otherwise}\end{cases}$
24:  end for
25:  Return extracted watermark $W$
IV Experimental Result
This section presents evaluation of the proposed framework. Section IV-A describes the datasets and preprocessing and augmentation strategies used for training and evaluation. Section IV-B analyzes the quality and stability of the learned invariant features, including feature consistency under distortion and the quality of image reconstruction. Section IV-C compares our method with existing self-supervised feature extractors, such as SimCLR, BYOL, and VICReg, using cosine similarity and linear evaluation metrics to assess feature robustness. Finally, Section IV-D demonstrates the effectiveness of our learning-based zero-watermarking scheme, providing comparisons with competitive baselines under a range of distortions.
IV-A Dataset and Feature Training
To evaluate the effectiveness and robustness of the proposed framework, we conducted experiments across multiple datasets, each selected to serve a specific purpose. All images were resized to a resolution of $128\times 128\times 3$ and normalized to the range $[0,1]$ to ensure consistency and stable convergence during training.
For invariant feature learning and analysis (Section IV-B), we used a subset of 40,000 images from the MSCOCO dataset, which contains a diverse range of real-world scenes. To assess generalization, we evaluated the trained models on 5,000 MSCOCO images, as well as 5,000 images each from the Imagenette and Flickr30k datasets. These datasets introduce unseen variations in content and background, helping to validate the stability of the learned representations under domain shifts. Additionally, high-resolution robustness was tested using 900 images from the DIV2K dataset.
For linear evaluation (Section IV-C), we used 10,000 labeled images from the CIFAR-100 dataset to train a downstream classifier on frozen features, and evaluated it on a separate test set of 3,000 images. CIFAR-100 provides a well-established benchmark for assessing semantic discriminability of features across object categories.
The feature extractor was trained using the Adam optimizer with separate learning rates for each component: $\eta_{\mathrm{FE}}=0.0005$ for the feature extractor, $\eta_{D}=0.001$ for the discriminator, and $\eta_{R}=0.0001$ for the reconstructor. As shown in Figure 5, the training losses of all three components converge stably over time: the discriminator and reconstructor losses decrease steadily, while the feature extractor’s adversarial loss increases as it successfully fools the discriminator, indicating effective minimax optimization and stable training dynamics.
Figure 5: Training Loss Curves for Invariant Feature Learning.
Since the proposed framework learns and generalizes invariance through input distortions, we apply a range of geometric and photometric transformations during training and testing. Table I summarizes the distortion strategies used. Notably, during training, only geometric distortions (e.g., rotation, translation, shear) are applied, while at test time, a broader spectrum of distortions—including blurring, compression, and noise—is introduced. This design enables us to evaluate the generalization capacity of the learned invariant features: although trained only with geometric distortions, the model demonstrates invariance to a wide variety of unseen image processing attacks such as salt-and-pepper noise and JPEG compression.
TABLE I: Summary of Distortions in Training and Testing.
Distortion Type Training Testing
Rotation $\pm$15° $\pm$25°
Width Shift 15% 25%
Height Shift 15% 25%
Shear 15% 25%
Zoom 15% 25%
Horizontal Flip Yes Yes
Gaussian Blur No Yes
Solarization No Yes
Crop / Cutout No Yes
JPEG Compression No Yes
Brightness Adjustment No Yes
Contrast Adjustment No Yes
Hue / Saturation Adjustment No Yes
Gaussian Noise No Yes
Salt & Pepper Noise No Yes
IV-B Analysis of the Proposed Invariant Feature
To assess the effectiveness of the proposed invariant feature learning framework, we conduct a detailed analysis of feature robustness and reconstruction quality across a variety of datasets and distortions. We divide the distortions into two categories: (1) common image processing distortions, such as blur, noise, and compression, and (2) geometric transformations, such as rotation, translation, and zoom. Our evaluation is performed on four datasets: MSCOCO, ImageNet, Flickr30k, and DIV2K, selected to cover a wide spectrum of image content and resolutions.
To quantify feature invariance, we compute the cosine similarity between the invariant feature vectors extracted from the original image $x$ and a distorted version $x^{\prime}$. A higher similarity indicates stronger invariance to the applied transformation. Quality of the the proposed reconstruction is evaluated using the peak signal-to-noise ratio (PSNR) between the original image $x$ and its reconstruction $\hat{x}$ from the extracted features. These two metrics together offer a comprehensive view of both the stability and informativeness of the learned representations.
Table II reports the cosine similarity and PSNR under common image processing distortions. Despite being trained only with geometric augmentations, the model demonstrates strong generalization to unseen distortions. For instance, cosine similarity remains consistently high under blur and JPEG compression ($\geq$0.93 across all datasets), indicating minimal impact on the feature representation. Gaussian noise and solarization result in modest drops in similarity, while salt-and-pepper noise introduces the largest variation, particularly on MSCOCO (0.92). Notably, Flickr30k and ImageNet show higher resilience to these pixel-level attacks, with cosine similarity values reaching up to 0.96.
The PSNR scores corroborate the feature analysis. MSCOCO, ImageNet, and Flickr30k achieve high reconstruction fidelity under most distortions, with blur and JPEG yielding PSNR values above 30 dB. In contrast, DIV2K—due to its higher resolution and finer detail—shows more degradation under extreme perturbations, particularly salt-and-pepper noise (27.3 dB) and solarization (26.0 dB). These results confirm that the learned features are both distortion-invariant and semantically expressive across multiple content domains.
TABLE II: Feature Cosine similarity and Image Reconstruction PSNR under common image processing distortions.
Dataset Cosine Similarity (Feature Invariance) $\uparrow$ PSNR (dB) $\uparrow$
Blur JPEG Gaussian Noise Salt & Pepper Solarization Blur (1.5) JPEG Gaussian Noise Salt & Pepper Solarization
MSCOCO 0.97 0.96 0.96 0.92 0.93 32.5 31.2 30.9 28.8 29.5
ImageNet 0.95 0.94 0.93 0.96 0.92 33.9 30.8 31.5 29.2 30.3
Flickr30k 0.98 0.96 0.95 0.96 0.95 32.2 31.8 29.8 26.5 30.2
DIV2K 0.95 0.93 0.93 0.91 0.93 29.5 31.0 28.7 27.3 26.0
Table III presents the results under geometric transformations. Rotation shows the highest feature invariance across all datasets (cosine similarity $\geq$0.95), followed by width and height shifts. Shear and zoom introduce slightly more disruption, with cosine similarity values dropping to 0.91–0.94 for ImageNet and DIV2K. Among geometric distortions, zoom exhibits the most pronounced impact on feature consistency, likely due to scale-dependent feature extraction challenges.
Reconstruction quality remains high for most geometric transformations, with PSNR values frequently exceeding 30 dB. For example, width shift on ImageNet achieves 33.5 dB, while even under rotation, MSCOCO retains a PSNR of 33.5 dB. However, zoom transformations reduce PSNR across all datasets, especially for Flickr30k (28.6 dB), indicating that scale variations affect both the spatial detail and feature structure more than other affine transformations.
TABLE III: Feature Cosine similarity and Image Reconstruction PSNR under geometric transformations.
Dataset Cosine Similarity (Feature Invariance) $\uparrow$ PSNR (dB) $\uparrow$
Rotation Width Shift Height Shift Shear Zoom Rotation Width Shift Height Shift Shear Zoom
MSCOCO 0.97 0.96 0.96 0.94 0.95 33.5 31.1 31.6 32.3 29.6
ImageNet 0.95 0.94 0.94 0.91 0.93 31.9 33.5 29.2 31.8 29.1
Flickr30k 0.97 0.95 0.94 0.94 0.92 30.2 33.7 31.7 30.2 28.6
DIV2K 0.96 0.92 0.93 0.92 0.94 31.5 32.1 30.8 32.8 29.6
Overall, the results demonstrate that the proposed framework effectively learns invariant features with strong generalization to unseen distortions, including both common image processing effects and geometric spatial transformations. To illustrate that the extracted features preserve not only invariance but also rich semantic information, we validate the reconstruction results on the testing data, and a few representative examples are shown in Figure 6.
Figure 6: Representative Examples of Image Reconstruction From Invariant Features.
IV-C Comparative Analysis of the Proposed Invariant Feature
To evaluate the robustness and effectiveness of the proposed invariant feature learning framework, we compare it with five representative self-supervised learning (SSL) baselines: SimCLR [17], BYOL [18], Barlow Twins [25], VICReg [26], and VIBCREG [27]. These methods were selected for their strong performance across SSL benchmarks and for covering diverse learning paradigms—including contrastive learning, redundancy reduction, and variance-invariance-covariance regularization. We assess the learned representations along two key axes: (1) feature invariance, i.e., the stability of feature embeddings under distortions, and (2) semantic descriptiveness, i.e., how well the features support classification when linearly probed. These two criteria together offer a comprehensive measure of robustness and utility for downstream watermarking tasks.
IV-C1 Invariant Feature Similarity
We first evaluate the degree to which each method preserves feature consistency under perturbations. Given an input image and a distorted variant, we extract feature embeddings using each method and compute the cosine similarity between the original and distorted features. A higher similarity indicates stronger invariance to the applied transformation. We test against five common distortions: Gaussian blur, JPEG compression, Gaussian noise, salt & pepper noise, and hue shift.
Table IV reports the results across four datasets. The proposed method consistently outperforms the SSL baselines across all distortions and datasets. For instance, under Gaussian blur and JPEG compression, our model achieves cosine similarities up to 0.98, demonstrating high resistance to low-level degradations. Even under challenging conditions like salt & pepper noise and hue shifts, our model maintains higher feature consistency than all compared methods. These results validate the effectiveness of our adversarial invariant learning objective in producing distortion-invariant feature embeddings.
TABLE IV: Quantitative Comparison of Feature Invariance (Cosine Similarity) across SSL Methods Under Diverse Distortions on Multiple Datasets.
Dataset Method Cosine Similarity (Feature Invariance) $\uparrow$
Gaussian Blur (1.5) JPEG (15) Gaussian Noise (0.15) Salt & Pepper (0.15) Hue (0.2)
MSCOCO SimCLR 0.97 0.92 0.92 0.94 0.90
VICReg 0.92 0.86 0.84 0.91 0.89
Barlow Twins 0.93 0.88 0.87 0.90 0.92
BYOL 0.96 0.93 0.91 0.93 0.95
VIbCReg 0.94 0.95 0.91 0.88 0.91
Ours 0.98 0.96 0.95 0.93 0.93
ImageNet SimCLR 0.94 0.87 0.89 0.92 0.90
VICReg 0.92 0.89 0.92 0.93 0.92
Barlow Twins 0.93 0.90 0.96 0.91 0.91
BYOL 0.94 0.91 0.92 0.95 0.89
VIbCReg 0.95 0.92 0.88 0.93 0.90
Ours 0.96 0.94 0.94 0.95 0.92
Flickr30k SimCLR 0.97 0.96 0.93 0.93 0.94
VICReg 0.95 0.93 0.95 0.92 0.92
Barlow Twins 0.94 0.93 0.94 0.94 0.93
BYOL 0.95 0.92 0.98 0.93 0.95
VIbCReg 0.96 0.93 0.94 0.94 0.91
Ours 0.98 0.96 0.94 0.95 0.95
DIV2K SimCLR 0.95 0.88 0.91 0.87 0.91
VICReg 0.94 0.92 0.90 0.93 0.92
Barlow Twins 0.94 0.91 0.92 0.87 0.92
BYOL 0.95 0.93 0.90 0.90 0.91
VIbCReg 0.93 0.90 0.91 0.91 0.90
Ours 0.95 0.93 0.94 0.92 0.93
IV-C2 Linear Evaluation for Semantic Discriminability
To assess whether the learned features also retain semantic richness, we conduct a linear evaluation protocol. For each SSL method, including ours, we extract features from a frozen encoder and train a linear classifier on 10,000 clean CIFAR-100 training images. The classifiers are then evaluated on distorted test sets from both CIFAR-100 and Imagenette, using the same five distortion types.
Figure 7: Linear Evaluation of Representation Quality: Top-1 and Top-5 Classification Accuracy Across SSL Methods on CIFAR-100 and Imagenette Under Various Distortions.
Figure 7 presents heatmaps of Top-1 and Top-5 linear classification accuracy across CIFAR-100 and Imagenette under five types of distortions. The proposed method consistently achieves the highest accuracy across both datasets and distortion types, outperforming all compared SSL baselines. For instance, under Gaussian noise, our method achieves a Top-1 accuracy of 75.8% on CIFAR-100 and 82.9% on Imagenette, significantly surpassing the next best performers. Similarly, in the Top-5 evaluation, our method yields 92.0% and 92.8% accuracy under noise on CIFAR-100 and Imagenette, respectively. These results highlight the effectiveness of our approach not only in learning invariant representations but also in preserving semantic discriminability across object classes, making it well-suited for downstream watermarking tasks that demand both robustness and generalization. A sample set of original and distorted images used in these experiments is illustrated in Figure 8.
Figure 8: Visual Illustration of Image Distortions used for Robustness Evaluation
IV-D Robust Zero Watermarking
This subsection presents an evaluation of the proposed learning-based zero-watermarking scheme using the invariant features extracted in Module 1. We organize the analysis into three parts. First, we visualize the training dynamics of the reference signature $C$, including t-SNE projections before and after optimization, and monitor the convergence behavior of the watermark loss over time. Next, we quantify the robustness of the watermarking system across a wide range of image distortions by evaluating bit accuracy under varying transformation intensities. Finally, we compare our method against existing state-of-the-art zero-watermarking and other deep watermarking approaches, highlighting the performance gains achieved by our framework in terms of robustness and stability. Together, these experiments establish the practicality and generalizability of our proposed method. We evaluated the proposed method using binary watermark messages of varying lengths and report results using a 30-bit watermark to align with common practice in the literature and ensure fair comparison with prior work.
IV-D1 Visualization of Reference Signature Learning
To understand the optimization dynamics of the learned reference signature $C$, we visualize both its geometric evolution in feature space and the corresponding training loss over time. Figure 9 presents two complementary views of this process: the structural arrangement of $C$ via t-SNE projection, and the loss curves from training the watermark for a few representative image-watermark pairs.
Figure 9: t-SNE Projection of Reference Signatures $C$ Before and After Training.
As shown in Fig. 9, we apply t-distributed Stochastic Neighbor Embedding (t-SNE) to project the high-dimensional reference vectors $C\in\mathbb{R}^{k\times d}$ into two dimensions, both before and after training. Each point represents the projection of one bit vector $C_{i}$ associated with a distorted image. Before training, the projected vectors appear scattered without clear structure, reflecting the random initialization. After optimization, the vectors form distinct and compact clusters in the embedding space. This emergent separation indicates that the learning process successfully aligns the reference signature with the distortion-invariant feature space. Notably, the improved spatial organization mirrors the expected behavior of our theoretical formulation, in which $C$ is optimized to serve as a directional codebook that maximizes the margin between binary watermark bits. The optimization implicitly identifies a hyperplane such that $\tilde{F}\cdot C_{i}\geq 0.5$ or $<0.5$, depending on the bit label, enabling robust watermark recovery through simple dot-product thresholding.
Figure 10: Training Loss for Random Watermarking Images.
Figure 10 shows the training loss curves over 50 epochs for five randomly selected images. The consistent and monotonic reduction in loss illustrates effective convergence of the reference optimization process, with the steep early drop reflecting rapid alignment with the underlying feature distribution. Together with the t-SNE results, this confirms that the learned signatures not only generalize across distorted views but are also geometrically structured in a way that supports discriminative watermark decoding. These findings reinforce the strength of learning-based watermark representation as a robust alternative to traditional hand-crafted correlation schemes.
IV-D2 Distortion Tolerance Range
Figure 11: BER (%) under Various Distortion Types.
Figure 11 illustrates the Bit Error Rate (BER) performance of the proposed watermarking system under six common example distortions: JPEG compression, Gaussian blur, cropping, Gaussian noise, salt & pepper noise, and hue adjustment. BER is computed as the ratio of incorrectly extracted bits to the total number of embedded bits, expressed as a percentage. All results are averaged over 1,000 test images, each containing a 30-bit binary watermark, resulting in a total of 30,000 evaluated bits per distortion setting.
For JPEG compression, BER consistently decreases as the quality factor improves, confirming strong robustness to compression artifacts. Gaussian blur exhibits a gradual BER increase with higher blur radii, while cropping and hue adjustment maintain relatively low BER values even at high distortion intensities. This indicates that the proposed method effectively preserves watermark integrity under spatial and color-shift perturbations. Noise-based distortions—especially Gaussian noise and salt & pepper noise—introduce relatively more significant degradation, with BER rising sharply under higher noise levels. Nonetheless, the absolute BER values remain low across all cases, ranging from 0.05% to a maximum of 8%, highlighting the overall robustness of the proposed method.
IV-D3 Comparison with Existing Watermarking Methods
To evaluate the robustness of the proposed watermarking framework, we compare it against several representative models across three categories: (1) deep learning-based watermarking schemes that embed information into the image, (2) feature extractor-based watermarking methods that leverage pre-trained SSL models, and (3) zero-watermarking approaches with deep models. Each comparison focuses on bit-level accuracy or BER under various common distortions.
For a fair comparison, all methods were evaluated using a fixed watermark length of 30 bits per image. Additionally, watermarking embedding was controlled to ensure consistency across methods. In particular, for embedded watermarking schemes, watermarking were applied such that the resulting marked images maintained a PSNR of approximately 40 dB. In contrast, the proposed method and other zero-watermarking baselines do not alter the original image, preserving its fidelity by design.
TABLE V: Bit Accuracy (%) Comparison with Watermarking Models that Embed Information.
Distortion [28] [29] Ours
Gaussian Noise (0.06) 74.6 95.6 97.8
Gaussian Noise (0.08) 67.7 93.5 96.7
Gaussian Noise (0.10) 63.2 89.5 95.3
Salt & Pepper (0.05) 99.1 95.7 95.5
Salt & Pepper (0.10) 93.1 85.0 95.1
Salt & Pepper (0.15) 83.4 77.1 93.9
Adjust Hue (0.2) 65.1 94.0 99.8
Adjust Hue (0.4) 34.0 70.7 99.3
Adjust Hue (0.6) 18.1 42.4 99.1
Adjust Saturation (5.0) 99.9 95.2 99.9
Adjust Saturation (10.0) 99.6 93.3 99.4
Adjust Saturation (15.0) 98.5 90.1 98.7
GIF (64) 87.1 97.6 99.5
GIF (32) 76.8 95.7 97.7
GIF (16) 65.0 91.7 93.9
Resize Width (0.9) 99.3 99.9 99.9
Resize Width (0.7) 85.3 88.4 97.3
Resize Width (0.5) 66.5 67.1 94.8
Average 76.5 86.8 96.5
Table V presents bit accuracy results under a wide range of distortions. We compare our method with HIDDEN [28], a widely used encoder-decoder watermarking scheme, and [29], which uses adversarial training to improve robustness. Our model consistently achieves the highest accuracy across most transformations, particularly under color-related distortions such as hue and saturation adjustment, where alternative methods degrade significantly. Moreover, our model shows stronger robustness under severe noise (e.g., Gaussian 0.10, Salt & Pepper 0.15), outperforming others by large margins. The average accuracy improvement of our model demonstrates superior generalization across both geometric and photometric distortions.
TABLE VI: Bit Accuracy (%) Comparison with Pretrained Feature-Based Watermarking Models.
Distortion [7] [8] Ours
Identity 1.00 1.00 1.00
Rotation (25) 0.27 1.00 1.00
Crop (0.5) 1.00 1.00 1.00
Crop (0.1) 0.02 0.98 1.00
Resize (0.7) 1.00 1.00 1.00
Blur (2.0) 0.25 1.00 1.00
JPEG (50) 0.96 0.97 0.99
Brightness (2.0) 0.99 0.96 1.00
Contrast (2.0) 1.00 1.00 1.00
Hue (0.25) 1.00 1.00 1.00
Screenshot 0.86 0.97 0.98
Table VI compares our framework with two recent methods that utilize pre-trained SSL models for watermark retrieval: Vukotić et al. [7], which uses convolutional networks, and Fernández et al. [8], which employs DINO features. Although these approaches demonstrate strong performance under identity-preserving transformations, they often struggle with heavy attacks and geometric alterations such as rotation or crop. In contrast, our model maintains perfect or near-perfect bit accuracy across all comparison distortions. Notably, under non-differentiable transformations like screenshots, our method preserves 98% bit accuracy, further validating the resilience conferred by our invariant feature learning.
TABLE VII: BER (%) Comparison across Zero-Watermarking Models.
Distortion [5] Ours
Gaussian Noise (0.005) 0.0151 0.0139
Gaussian Noise (0.01) 0.0231 0.0187
JPEG Compression (30) 0.0267 2.0314
JPEG Compression (70) 0.0169 0.5748
Rotation (3) 0.0473 0.0196
Rotation (5) 0.0622 0.0237
Scaling (0.5) 0.0182 0.0143
Scaling (2.0) 0.0153 0.0093
Median Filter ($3\times 3$) 0.0127 0.0151
Median Filter ($5\times 5$) 0.0170 0.0291
Gaussian Filter ($3\times 3$) 0.0075 0.0038
Gaussian Filter ($5\times 5$) 0.0126 0.0089
Table VII reports BER comparisons with the recent deep zero-watermarking method from [2]. Across different distortion types—including compression, filtering, and geometric transformations—our method demonstrates competitive or superior robustness. In particular, under geometric distortions such as rotation and scaling, and under common image processing operations like Gaussian filtering, our framework achieves consistently lower BER, with improvements up to 50% in some cases (e.g., scaling at factor 2.0 and Gaussian filter $3\times 3$). While our method shows slightly higher BER under extreme JPEG compression, it maintains strong robustness across a broader range of transformations. These results underscore the generalizability and robustness of the learned invariant features for zero-watermarking under diverse perturbations.
These results collectively demonstrate the superiority of our approach across different methodological paradigms. Unlike existing approaches that rely on embedded information or static feature extractors, our method learns an adaptive and robust signature in an invariant feature space, offering strong performance under a broad range of real-world transformations.
V Conclusion
This paper presents InvZW, a deep learning-based framework for robust image zero-watermarking built upon invariant feature learning through noise-adversarial training. Unlike traditional or pretrained feature extractor-based methods, our approach explicitly learns distortion-invariant yet semantically meaningful representations via a combination of adversarial alignment and reconstruction constraints. These representations serve as a stable basis for a novel zero-watermarking scheme, where a binary watermark is associated with a reference signature optimized in the invariant feature space.
We demonstrate that the proposed framework not only achieves high robustness against diverse photometric and geometric distortions, but also enables accurate watermark recovery through a novel flexible embedding mechanism. Extensive evaluations show superior performance over representative deep watermarking methods, including encoder-decoder-based, self-supervised, and zero-watermarking baselines.
Future work will focus on extending the framework to video watermarking and real-time applications, as well as incorporating adaptive optimization strategies for signature learning under dynamic distortion environments. The integration of learned invariance and adaptive encoding opens new possibilities for secure, high-fidelity media protection in both static and streaming contexts.